# Seek-CAD: A Self-refined Generative Modeling for 3D Parametric CAD Using Local Inference via DeepSeek

### v2

@@ Abstract
The advent of Computer-Aided Design (CAD) generative modeling will significantly transform the design of industrial products. The recent research endeavor has extended into the realm of Large Language Models (LLMs). In contrast to fine-tuning methods, training-free approaches typically utilize the advanced closed-source LLMs, thereby offering enhanced flexibility and efficiency in the development of AI agents for generating CAD parametric models. However, the substantial cost and limitations of local deployment of the top-tier closed-source LLMs pose challenges in practical applications. The Seek-CAD is the pioneer exploration of locally deployed open-source inference LLM DeepSeek-R1 for CAD parametric model generation with a training-free methodology. This study is the first investigation to incorporate both visual and Chain-of-Thought (CoT) feedback within the self-refinement mechanism for generating CAD models. Specifically, the initial generated parametric CAD model is rendered into a sequence of step-wise perspective images, which are subsequently processed by a Vision Language Model (VLM) alongside the corresponding CoTs derived from DeepSeek-R1 to assess the CAD model generation. Then, the feedback is utilized by DeepSeek-R1 to refine the initial generated model for the next round of generation. Moreover, we present an innovative 3D CAD model dataset structured around the SSR (Sketch, Sketch-based feature, and Refinements) triple design paradigm. This dataset encompasses a wide range of CAD commands, thereby aligning effectively with industrial application requirements and proving suitable for the generation of LLMs. Extensive experiments validate the effectiveness of Seek-CAD under various metrics.
@@ 1 Introduction
The CAD parametric model (also called design history), as a crucial role in Computer-Aided Design (CAD), indicates the design logic of 3D CAD models, where its command and parameter sequences can be quickly edited to create or modify the shape of a 3D object  [54, 21, 32]. However, constructing a parametric CAD model from scratch is time-consuming and hinders the development of automation in industrial manufacturing. Hence, it attracts much attention on generative CAD modeling recently, including many interesting applications such as CAD parts assembly  [20, 52, 53], shape parsing  [24, 40], CAD parametric model generation  [54, 57, 29], and cross-modal CAD generation (e.g., Point cloud-to-CAD, Text-to-CAD, and etc.) [22, 26, 8, 28, 27, 33, 30, 63, 23]. Under the current trend of LLMs and VLMs performing outstandingly in computer vision tasks [50, 41, 47, 16, 64, 3], integrating LLMs and VLMs in generative CAD modeling will pave the way for future innovations in smart design systems.
@@
Fine-tuning is a commonly used method for adapting general LLMs to domain-specific applications. In contrast to fine-tuning approaches, training-free approaches typically utilize advanced closed-source LLMs i.e. GPT-4o [1], thereby offering enhanced flexibility and efficiency in the creation of AI agents for generating CAD parametric models [61, 2, 25]. However, the substantial cost and limitations of the local deployment of the top-tier closed-source LLMs pose challenges in practical applications. Inspired by recent DeepSeek-R1 [12] with advanced capability of reasoning, we deploy a DeepSeek-R1-32B-Q4 locally without training or finetuning as the backbone of our approach to explore its capability for generative CAD modeling. To the best of our knowledge, this is the first attempt to explore DeepSeek-R1 on generative CAD modeling.
@@
Specifically, we present Seek-CAD, a training-free generative framework for CAD modeling based on DeepSeek-R1-32B-Q4. By employing a retrieval-augmented generation (RAG) strategy on a local CAD code corpus, Seek-CAD produces Python-like CAD code for CAD parametric model generation. To ensure that the generated code aligns with prompt descriptions and faithfully encodes geometric features (e.g., fillet, chamfer) and constraints (e.g., tangency, orthogonality), we introduce a step-wise visual feedback mechanism that guides and refines the modeling process. In particular, we render step-wise perspective images that visually capture each stage of the CAD modeling process. These images are then evaluated using Gemini-2.0 [45] to assess their alignment with the chain-of-thought (CoT) from DeepSeek-R1-32B-Q4. The resulting feedback is incorporated to iteratively refine the initial code and parameters.
@@
Furthermore, we propose a novel CAD design paradigm called SSR (Sketch, Sketch-based feature, and Refinements), where each model is represented as a sequence of SSR triples, each consisting of a sketch, a sketch-based feature (e.g., extrude and revolve), and optionally, refinement features (e.g., chamfer, fillet and shell). Complex shapes are constructed through boolean operations across SSR units. To support refinement features, we introduce a simple yet effective reference mechanism, termed CapType (Figure 3), which establishes explicit links between topological primitives in the sketch and their resulting primitives generated during modeling. To evaluate our approach, we construct a new CAD dataset of 40k samples following the SSR modeling paradigm. The dataset covers diverse CAD features not included in existing datasets, and each sample is paired with a textual description generated by GPT-4o [1]. For further details, please see Section A.2. We plan to release the dataset publicly in the future.
@@ 2.1 Generative CAD Modeling
Generative CAD modeling has made a significant step recently [56, 58, 18, 8, 13], notably with transformer-based models, which treat CAD commands as parametric sequential data for learning and generation with a feed-forward or an auto-regressive strategy [54, 56, 57, 10]. Besides, diffusion-based [15, 33] methods have also been adopted to achieve the controllable generation or reconstruction of parametric CAD sequences [33, 62]. Mamba-CAD [29] makes a step forward to handle longer parametric CAD sequences to generate complex CAD models. As a CAD model contains explicit parametric commands, which implicitly indicate its design logic and shape geometry, it can be seen as a kind of multi-modal data inherently, which also raises many applications of generating or reconstructing CAD models from point clouds, images, and texts [27, 60, 8, 22, 28]. These methods mainly focus on the SE (Sketch-Extrusion) paradigm, which supports only a limited set of simple CAD operations, making it incapable of generating diverse and complex CAD models that meet real-world design requirements. Instead, Seek-CAD adopts the SSR paradigm, which enables the inclusion of diverse CAD commands such as fillet, chamfer, and shell, supporting the creation of more complex CAD models, which is closer to industrial requirements. From another perspective, Seek-CAD is a self-refined framework for generative CAD modeling without any training or finetuning, which is also different from these efforts under a training strategy.
@@ 2.2 LLMs For CAD
With the success of LLMs in Computer Vision (CV) tasks [44, 35, 42, 11], many interesting applications of leveraging LLMs for CAD-related tasks have raised much attention recently.Text2CAD [23] generates parametric CAD models from natural language instructions using a transformer-based network [46] with an annotation pipeline for CAD prompt using Mistral [4] and LLaVA-NeXT [31]. CAD-MLLM [55], a unified system to utilize a LLM to align multimodal inputs with parametric CAD sequences for generating CAD models. CAD-assistant [34] uses a Vision-Large Language Model (VLLM) with tool-augmented planning to iteratively generate and adapt CAD designs via Python API of [9]. FI2CAD [36] introduces a multi-agent system, a LLM-based MAS architecture for CAD development processes that mimics an engineering team to automatically generate and refine parametric CAD models with human feedback. CAD-Llama [25] enables pretrained LLMs to generate parametric 3D CAD models through the adaptive pretraining on Structured Parametric CAD Code (SPCC). Unlike these methods, our approach mainly focuses on a training-free strategy with the self-refined capability for generative CAD modeling.
The most related to our Seek-CAD are other two training-free frameworks, 3D-PreMise [61] and CADCodeVerify [2]. 3D-Premise enhances CAD code refinement by supplying GPT-4 [1] with an image of a whole object and its initial description, allowing it to identify and correct discrepancies between the intended design and the generated CAD code. Unlike 3D-Premise, CADCodeVerify prompts GPT-4 [1] to generate and answer a set of questions based on the initially provided description of a 3D CAD object, to adjust any discrepancies between the generated CAD model and the description. Compared to them, Seek-CAD differs in the following aspects: (i) Seek-CAD uses DeepSeek-R1-32B-Q4, an open-source reasoning model for local deployment, integrated with a CAD code corpus via a retrieval-augmented generation framework, instead of relying on proprietary LLMs like GPT-4o. (ii) Previous research evaluates object renderings in VLMs by focusing only on final CAD forms, overlooking intermediate phases, which limits feedback for complex models. Seek-CAD addresses this with a step-wise feedback mechanism, showing both final and intermediate shapes to enhance VLM feedback. (iii) Previous methods use predefined question templates to evaluate VLM alignment with descriptions and images. Seek-CAD, however, guides VLMs to assess alignment between DeepSeek-R1’s chain-of-thought (CoT) and step-wise images, as CoT effectively illustrates the design logic, enabling clearer VLM understanding (Sec 3.2).
@@ 3 Seek-CAD Framework
Figure 1: The overview of our Seek-CAD framework. The whole pipeline can be divided into two parts consisting of "Initial CAD Code Generation" and "CAD Code Refinement", which are both embedded with a knowledge constraint depicted in Sec. 3.1 to guide DeepSeek-R1 to generate CAD code following the SSR paradigm (Sec. 4). For the first part, a given query $T$ is enhanced by conducting RAG on a local CAD corpus that consisting 10, 000 CAD models. Next, Top-3 retrieved candidates would be concatenated with $T$ to trigger DeepSeek-R1 to generate an initial CAD code $I_{0}$. For the second part, $I_{0}$ would go through the Step-wise Visual Feedback with CoT to have the iteration refinement. To achieve this, we first utilize a rendering script $R(*)$ to obtain step-wise images of $I_{0}$, which can represents the intermediate and ultimate shape of the object ($M_{I}$, $M_{U}$) simultaneously. $\oplus$ denotes the concatenation of SSR triplets, where each triplet, represented by $S_{i}$ (Sec. 4), is rendered along with all its preceding triplets to preserve the correlations between object entities. (More details in Sec. 3.2). Next, the step-wise images are fed into Gemini-2.0 to assess their alignment with the CoT from DeepSeek-R1. This feedback determines whether the current code $I_{k}$ is reasonable. In practice, we set $k=2$ as the maximize iterations of code refinement.
@@ 3.1 Local Inference Pipeline
(2) Knowledge Constraint.  Similar to other LLMs, DeepSeek-R1 may exhibit hallucinations [43, 17, 19], occasionally generating CAD parametric models that deviate from the SSR paradigm. To address it, we propose the knowledge constraint $Cons=(\Phi,\mathcal{D},\mathcal{E})$ as the system prompt to make DeepSeek-R1 generate CAD code following the SSR paradigm, with the constraint $Cons$ consisting of three parts: $\Phi$ specifies its functionality, $\mathcal{D}$ documents the SSR schema, and $\mathcal{E}$ provides an example pairing a textual description with SSR-based CAD code, as shown in Figure 6 of Appendix A.1.
(3) Retrieval Augment Generation (RAG).  In order to augment the recognition of Seek-CAD within the SSR paradigm, we construct a SSR-based CAD corpus $\mathcal{C}_{\text{SSR}}=\{(d_{i},c_{i})\}_{i=1}^{N}$ to equip Seek-CAD with the capability to perform RAG. $d_{i}$ is the textual description and $c_{i}$ its corresponding CAD code, with $N=10,\!000$ CAD models. As usual settings of conducting RAG, we select a hybrid search combining a vector-based and full-text strategy. Specifically, for a given $T$, the similarity score is computed by:
@@
where $g_{i}^{\text{vec}}=\cos(z_{T},z_{i}),\quad z_{T}=\texttt{e}_{\text{vec}}(T),\quad z_{i}=\texttt{e}_{\text{vec}}(d_{i})$. $\texttt{e}_{\text{vec}}$ denotes an embedding model, bge-m3 [5]. $g_{i}^{\text{full}}=\texttt{U}_{index}(T,d_{i})$ is calculated with a traditional inverted index without vectorization. The top-$k$ candidates are selected as $\mathcal{R}_{T}=\texttt{TopK}(g_{i}^{\text{final}})$. Practically, we set $\lambda=0.3$, top-$k=3$.
@@
The initially generated CAD code $I_{0}$ is occasionally subject to compilation failures with the geometry kernel (e.g., PythonOCC [39] utilized in this study), which are attributed to syntax errors E. For remediation, a pattern template Q is employed to automatically rectify E in $I_{0}$, addressing issues such as mismatched parentheses and incorrect capitalization of variable names, depicted as: $I_{0}\sim P(I_{0}\mid Q)$.

### v3

# arXiv:2505.17702v3[cs.CV] 28 Feb 2026

@@ Abstract
The advent of Computer-Aided Design (CAD) generative modeling will significantly transform the design of industrial products. The recent research endeavor has extended into the realm of Large Language Models (LLMs). In contrast to fine-tuning methods, training-free approaches typically utilize the advanced LLMs, thereby offering enhanced flexibility and efficiency in the development of AI agents for generating CAD parametric models. However, the lack of a mechanism to harness Chain-of-Thought (CoT) limits the potential of LLMs in CAD applications. The Seek-CAD is the pioneer exploration of locally deployed inference LLM DeepSeek-R1 for CAD parametric model generation with a training-free methodology. This study is the investigation to incorporate both visual and CoT feedback within the self-refinement mechanism for generating CAD models. Specifically, the initial generated parametric CAD model is rendered into a sequence of step-wise perspective images, which are subsequently processed by a Vision Language Model (VLM) alongside the corresponding CoTs derived from DeepSeek-R1 to assess the CAD model generation. Then, the feedback is utilized by DeepSeek-R1 to refine the initial generated model for the next round of generation. Moreover, we present an innovative 3D CAD model dataset structured around the SSR (Sketch, Sketch-based feature, and Refinements) triple design paradigm. This dataset encompasses a wide range of CAD commands, thereby aligning effectively with industrial application requirements and proving suitable for the generation of LLMs. Extensive experiments validate the effectiveness of Seek-CAD under various metrics.
@@ 1 Introduction
The CAD parametric model (also called design history), as a crucial role in Computer-Aided Design (CAD), indicates the design logic of 3D CAD models, where its command and parameter sequences can be quickly edited to create or modify the shape of a 3D object  Wu et al. (2021), Jones et al. (2023). However, constructing a parametric CAD model from scratch is time-consuming and hinders the development of automation in industrial manufacturing. Hence, it attracts much attention on generative CAD modeling recently, including many interesting applications such as CAD parts assembly  Jones et al. (2021), Willis et al. (2022), Wu et al. (2023), shape parsing  Li et al. (2022), Ren et al. (2022), CAD parametric model generation  Wu et al. (2021), Xu et al. (2023), Li et al. (2025c), and cross-modal CAD generation (e.g., Point cloud-to-CAD, Text-to-CAD, and etc.) Khan et al. (2024a), Li et al. (2024a), Dupont et al. (2024), Li et al. (2024c; b), Ma et al. (2024), Li et al. (2025d), Zhang et al. (2025b), Khan et al. (2024b). Under the current trend of LLMs and VLMs performing outstandingly in computer vision tasks Wang et al. (2023), Schumann et al. (2024), Wang and Ke (2024), Huang et al. (2024), Zhu et al. (2024), Bensabath et al. (2024), integrating LLMs and VLMs in generative CAD modeling will pave the way for future innovations in smart design systems.
@@
Fine-tuning is a commonly used method for adapting general LLMs to domain-specific applications. In contrast to fine-tuning approaches, training-free approaches typically utilize advanced LLMs i.e. GPT-4o Achiam et al. (2023), thereby offering enhanced flexibility and efficiency in the creation of AI agents for generating CAD parametric models Yuan et al. (2024), Alrashedy et al. (2025). However, the primary drawback is the absence of a mechanism to harness Chain-of-Thought (CoT), which limits the potential of LLMs in CAD applications. Inspired by DeepSeek-R1 Guo et al. (2025) with advanced capability of reasoning, we deploy a DeepSeek-R1-32B-Q4 locally without training or finetuning as the backbone of our approach to explore its capability for generative CAD modeling.
@@
Specifically, we present Seek-CAD, a training-free generative framework for CAD modeling based on DeepSeek-R1-32B-Q4. By employing a retrieval-augmented generation (RAG) strategy on a local CAD code corpus, Seek-CAD produces Python-like CAD code for CAD parametric model generation. To ensure that the generated code aligns with prompt descriptions and faithfully encodes geometric features (e.g., fillet, chamfer) and constraints (e.g., tangency, orthogonality), we introduce a step-wise visual feedback mechanism that guides and refines the modeling process. In particular, we render step-wise perspective images that visually capture each stage of the CAD modeling process. These images are then evaluated using Gemini-2.0 Team et al. (2024) to assess their alignment with the chain-of-thought (CoT) from DeepSeek-R1-32B-Q4. The resulting feedback is incorporated to iteratively refine the initial code and parameters.
@@
Furthermore, we propose a novel CAD design paradigm called SSR (Sketch, Sketch-based feature, and Refinements), where each model is represented as a sequence of SSR triples, each consisting of a sketch, a sketch-based feature (e.g., extrude and revolve), and optionally, refinement features (e.g., chamfer, fillet and shell). Complex shapes are constructed through boolean operations across SSR units. To support refinement features, we introduce a simple yet effective reference mechanism, termed CapType (Figure 3), which establishes explicit links between topological primitives in the sketch and their resulting primitives generated during modeling. To evaluate our approach, we construct a new CAD dataset of 40k samples following the SSR modeling paradigm. The dataset covers diverse CAD features not included in existing datasets, and each sample is paired with a textual description generated by GPT-4o Achiam et al. (2023). For further details, please see Section A.2. The datatset has been released publicly and can be acessed in https://github.com/Sunny-Hack/Seek-CAD.
@@ 2.1 Generative CAD Modeling
Generative CAD modeling has made a significant step recently Xu et al. (2022; 2024b), Jayaraman et al. (2021), Dupont et al. (2024), Guo et al. (2022), notably with transformer-based models, which treat CAD commands as parametric sequential data for learning and generation with a feed-forward or an auto-regressive strategy Wu et al. (2021), Xu et al. (2022; 2023), Ganin et al. (2021). Besides, diffusion-based Ho et al. (2020) methods have also been adopted to achieve the controllable generation or reconstruction of parametric CAD sequences Ma et al. (2024), Zhang et al. (2025a). Meanwhile, reinforcement learning offers another effective paradigm for generating CAD models Li et al. (2025a). Mamba-CAD Li et al. (2025c) makes a step forward to handle longer parametric CAD sequences to generate complex CAD models. As a CAD model contains explicit parametric commands, which implicitly indicate its design logic and shape geometry, it can be seen as a kind of multi-modal data inherently, which also raises many applications of generating or reconstructing CAD models from point clouds, images, and texts Li et al. (2024b), You et al. (2024), Dupont et al. (2024), Khan et al. (2024a), Li et al. (2024c). These methods mainly focus on the SE (Sketch-Extrusion) paradigm, which supports only a limited set of simple CAD operations, making it incapable of generating diverse and complex CAD models that meet real-world design requirements. Instead, Seek-CAD adopts the SSR paradigm, which enables the inclusion of diverse CAD commands (similar to the complex feature set in WHUCAD Fan et al. (2025)) such as fillet, chamfer, and shell, supporting the creation of more complex CAD models, which is closer to industrial requirements, which is closer to industrial requirements. From another perspective, Seek-CAD is a self-refined framework for generative CAD modeling without any training or finetuning, which is also different from these efforts under a training strategy.
@@ 2.2 LLMs For CAD
With the success of LLMs in Computer Vision (CV) tasks Tan et al. (2024), Naeem et al. (2023), Shao et al. (2024), Ghosh et al. (2024), many interesting applications of leveraging LLMs for CAD-related tasks have raised much attention recently.Text2CAD Khan et al. (2024b) generates parametric CAD models from natural language instructions using a transformer-based network Vaswani et al. (2017) with an annotation pipeline for CAD prompt using Mistral Chaplot (2023) and LLaVA-NeXT Liu et al. (2024). CAD-MLLM Xu et al. (2024a), a unified system to utilize a LLM to align multimodal inputs with parametric CAD sequences for generating CAD models. CAD-assistant Mallis et al. (2024) uses a Vision-Large Language Model (VLLM) with tool-augmented planning to iteratively generate and adapt CAD designs via Python API of FreeCAD . FI2CAD Ocker et al. (2025) introduces a multi-agent system, a LLM-based MAS architecture for CAD development processes that mimics an engineering team to automatically generate and refine parametric CAD models with human feedback. CAD-Llama Li et al. (2025b) enables pretrained LLMs to generate parametric 3D CAD models through the adaptive pretraining on Structured Parametric CAD Code (SPCC). Unlike these methods, our approach mainly focuses on a training-free strategy with the self-refined capability for generative CAD modeling.
The most related to our Seek-CAD are other two training-free frameworks, 3D-PreMise Yuan et al. (2024) and CADCodeVerify Alrashedy et al. (2025). 3D-Premise enhances CAD code refinement by supplying GPT-4 Achiam et al. (2023) with an image of a whole object and its initial description, allowing it to identify and correct discrepancies between the intended design and the generated CAD code. Unlike 3D-Premise, CADCodeVerify prompts GPT-4 Achiam et al. (2023) to generate and answer a set of questions based on the initially provided description of a 3D CAD object, to adjust any discrepancies between the generated CAD model and the description. Compared to them, Seek-CAD differs in the following aspects: (i) Seek-CAD uses DeepSeek-R1-32B-Q4 (open-source reasoning model) for local deployment, integrated with a CAD code corpus via a retrieval-augmented generation framework, instead of relying on GPT-4. (ii) Previous research evaluates object renderings in VLMs by focusing only on final CAD forms, overlooking intermediate phases, which limits feedback for complex models. Seek-CAD addresses this with a step-wise feedback mechanism, showing both final and intermediate shapes to enhance VLM feedback. (iii) Previous methods use predefined question templates to evaluate VLM alignment with descriptions and images. Seek-CAD, however, guides VLMs to assess alignment between DeepSeek-R1’s chain-of-thought (CoT) and step-wise images, as CoT effectively illustrates the design logic, enabling clearer VLM understanding (Sec 3.2).
@@ 3 Seek-CAD Framework
Figure 1: The overview of our Seek-CAD framework. The whole pipeline can be divided into two parts consisting of "Initial CAD Code Generation" and "CAD Code Refinement", which are both embedded with a knowledge constraint depicted in Sec. 3.1 to guide DeepSeek-R1 to generate CAD code following the SSR paradigm (Sec. 4). For the first part, a given query $T$ is enhanced by conducting RAG on a local CAD corpus that consisting 10, 000 CAD models. Next, Top-3 retrieved candidates would be concatenated with $T$ to trigger DeepSeek-R1 to generate an initial CAD code $I_{0}$. For the second part, $I_{0}$ would go through the Step-wise Visual Feedback with CoT to have the iteration refinement. To achieve this, we first utilize a rendering script R()R(*) to obtain step-wise images of $I_{0}$, which can represents the intermediate and ultimate shape of the object ($M_{I}$, $M_{U}$) simultaneously. $\oplus$ denotes the concatenation of SSR triplets, where each triplet, represented by $S_{i}$ (Sec. 4), is rendered along with all its preceding triplets to preserve the correlations between object entities. (More details in Sec. 3.2). Next, the step-wise images are fed into Gemini-2.0 to assess their alignment with the CoT from DeepSeek-R1. This feedback determines whether the current code $I_{k}$ is reasonable. In practice, we set $k=1$ as the maximize iterations of code refinement.
@@ 3.1 Local Inference Pipeline
(2) Knowledge Constraint. Similar to other LLMs, DeepSeek-R1 may exhibit hallucinations Sriramanan et al. (2024), Huang et al. (2025), Jiang et al. (2024), occasionally generating CAD parametric models that deviate from the SSR paradigm. To address it, we propose the knowledge constraint $Cons=(\Phi,\mathcal{D},\mathcal{E})$ as the system prompt to make DeepSeek-R1 generate CAD code following the SSR paradigm, with the constraint $Cons$ consisting of three parts: $\Phi$ specifies its functionality, $\mathcal{D}$ documents the SSR schema, and $\mathcal{E}$ provides an example pairing a textual description with SSR-based CAD code, as shown in Figure 6 of Appendix A.1.
(3) Retrieval Augment Generation (RAG). In order to augment the recognition of Seek-CAD within the SSR paradigm, we construct a SSR-based CAD corpus $\mathcal{C}_{\text{SSR}}=\{(d_{i},c_{i})\}_{i=1}^{N}$ to equip Seek-CAD with the capability to perform RAG. $d_{i}$ is the textual description and $c_{i}$ its corresponding CAD code, with N=10, [^0] N=10,\!000 CAD models. As usual settings of conducting RAG, we select a hybrid search combining a vector-based and full-text strategy. Specifically, for a given $T$, the similarity score is computed by:
@@
where $g_{i}^{\text{vec}}=\cos(z_{T},z_{i}),\quad z_{T}=\texttt{e}_{\text{vec}}(T),\quad z_{i}=\texttt{e}_{\text{vec}}(d_{i})$. $\texttt{e}_{\text{vec}}$ denotes an embedding model, bge-m3 Chen et al. (2024). $g_{i}^{\text{full}}=\texttt{U}_{index}(T,d_{i})$ is calculated with a traditional inverted index without vectorization. The top-$k$ candidates are selected as $\mathcal{R}_{T}=\texttt{TopK}(g_{i}^{\text{final}})$. Practically, we set $\lambda=0.3$, top-$k=3$.
@@
The initially generated CAD code $I_{0}$ is occasionally subject to compilation failures with the geometry kernel (e.g., PythonOCC Paviot (2022) utilized in this study), which are attributed to syntax errors E. For remediation, a pattern template Q is employed to automatically rectify E in $I_{0}$, addressing issues such as mismatched parentheses and incorrect capitalization of variable names, depicted as: $I_{0}\sim P(I_{0}\mid Q)$.